how many non-negative integers less than 104 that contain the digit 2 ?


1
Expert's answer
2022-02-07T10:21:50-0500

We identify one-digit number "a" with "000a," two-digit number "n=\overline{ab}" with "00ab," and three-digit number "n=\overline{abc}" with "0abc." Taking into account that there are ten digits, we conclude by Combinatorial product rule that the number of non-negative integers less than "10^4" is equal to "10\cdot 10\cdot 10\cdot 10=10,000." Since without digit 2 there are nine digits, we conclude by Combinatorial product rule that the number of non-negative integers less than "10^4" that do not contain the digit 2 is equal to "9\cdot 9\cdot 9\cdot 9=6,561."

We conclude that the number of non-negative integers less than 104 that contain the digit 2 is equal to "10,000-6,561=3,439."